Light is traveling from air to an ‘unknown material’. The angle of incidence is measured at 30.0o, and the angle of refraction is measured at 26.3o. What is the speed of light in this ‘unknown material’.
1
Expert's answer
2020-12-03T10:59:01-0500

Since we know that the angle of refraction is greater in media where the speed of light is greater, we can say that the speed of light in this material is smaller than in air:


"n=\frac{v_i}{v_r}=\frac{\text{sin}\theta_i}{\text{sin}\theta_r},\\\space\\
v_r=v_i\frac{\text{sin}\theta_r}{\text{sin}\theta_i}=3\cdot10^8\frac{\text{sin}26.3°}{\text{sin}30°}=265.8\cdot10^6\text{ m/s}."